ok please help me,
aead is 5 years older than Jen. Four times Jen's age increased by 3 years equals three times Ed's age diminished by 2 years. Find each person's age.
ed?
Jen?

Found 2 solutions by Fermat, jb007mi6:
Answer by Fermat(136)

Let E represent Ed's age
Let J represent Jen's age
The first part of the word problem says,
Ed is 5 years older than Jen: this becomes,
E = J + 5
==========
The second part of the word problem says,
Four times Jen's age increased by 3 years equals three times Ed's age diminished by 2 years: and this becomes,
4xJ + 3 = 3xE - 2
4J + 3 = 3E - 2
4J + 5 = 3E
now we substitute for E = J + 5 into the above eqn.
4J + 5 = 3(J + 5)
4J + 5 = 3J + 15
J = 10
======
E = J + 5
E = 15
======

Answer by jb007mi6(53)

Ed-5=Jen
4(Jen)+3=3(Ed)-2
use "E" and "J" from now on...
so if J=E-5, then you can plug "e-5" everywhere you see a J in the other equation
4(e-5)=3(e)-2
4e-20=3e-2
e=18
Ed is 18 yrs old
Jen is 13 yrs old